what is

20t=5t^2

Answer by Osran_Shri(18)

20t=5t^2
Rewrite the equation as :
5t^2 = 20 t
5t^2-20t = 0
5t ( t-4 ) = 0
By factors method:
5t = 0 and t-4 = 0
therefore t = 0, and 4
Here if you substitute value of t=0 then equation becomes undefined therefore
t = 4
Another simple Method:
5t^2 = 20 t
5t^2 / 20t = 1
t/4 = 1
Therefore t = 4 This is the simplest solution
Contact me for your assignments:
osran_con@yahoo.co.in